OpenStudy (anonymous):

One bulldozer exerts a force of 1850 pounds in a westerly direction. The other bulldozer pushes the piling with a force of 3250 pounds in a northerly direction. What is the magnitude of the resultant force upon the piling, to the nearest ten pounds?

OpenStudy (anonymous):

Welcome to OS! :)

OpenStudy (anonymous):

Draw a free body diagram, and using Pythagoras' theorem, You'll see that the total force is sqrt(1850^2+3250^2)

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

\[r^2 = 1850^2 + 3250^2 = 13985000 \rightarrow r = \sqrt{13985000} \approx 3740 pounds\]